OpenStudy (anonymous):

3x = 2y − 16 4x − y = -13

OpenStudy (anonymous):

x = -2 and y = 5

OpenStudy (callisto):

3x = 2y − 16 => 3x - 2y = -16 -(1) 4x − y = -13 => 4x − y = -13 -(2) (2)x2 - (1) 2(4x − y) - (3x - 2y) = 2(-13) - (-16 ) Solve x. Then put the value of x into (1) and solve y.

OpenStudy (anonymous):

Find value of y from second equation: \[y = 4x + 13\] Now put this in first equation: \[3x - 2(4x + 13) = -16 \implies 3x - 8x = -16 + 26 \implies -5x = 10\] Divide both the sides by -5 and get x and them by plugging value of x in any of the equation find y..

OpenStudy (anonymous):

3x-2y=-16.............1 4x-y=-13..............2